OpenStudy (mylahb):

Solving rational equations. Solve each equation and find LCD 6/x-1+2x/x-2=2

OpenStudy (anonymous):

First, you've gotta multiply the two denominators together to be able to add the numerators. So (x-1)(x-2), the fraction will end up looking like this: \[\frac{ 6+2x }{ x ^{2}-3x+2 }=2\] Your LCD (least common denominator) is x^2-3x+2 Do you know what to do next to solve? :)

OpenStudy (mylahb):

No

OpenStudy (anonymous):

Okiedoke, you wanna get x by itself, so the first thing you'd do is multiply both sides of the equation by the denominator, because it will cancel out on the left and leave this: 6+2x=2(x^2-3x+2) Distribute.. 6+2x=2x^2-6x+4 Combine like terms.. 0=2x^2-8x-2 At this point you use the quadratic formula to solve \[\frac{ -b + or - \sqrt{b ^{2}-4ac} }{ 2a }\] Your a=2, b=-8, c=-2 So plug those in and it'll look like: \[\frac{ 8 + or - \sqrt{64+16} }{ 4 }\] Simplify and it'll be: \[\frac{ 8 + or - \sqrt{80} }{ 4 }\] The sqrt80 can be divided to be sqrt16 times sqrt5, so it'll simplify to 4sqrt5 and look like: \[\frac{ 8 + or - 4\sqrt{5} }{ 4 }\] Now you can divide everything up top by the 4 below and you're left with: \[2+ or - \sqrt{5}\] So the solutions are 2+sqrt5 and 2-sqrt5 Hope that helps!!